A square and a rectangle each has a perimeter of 48 m. The difference between the areas of two figures is 4 m2. Find the dimensions of the rectangle.
Answer by math-vortex(648)

Hi, there--

THE PROBLEM:
A square and a rectangle each has a perimeter of 48 m. The difference between the areas of 
two figures is 4 square meters. Find the dimensions of the rectangle.

A SOLUTION:
Consider the square with perimeter 48m. 
Let x be the length of each side.

The perimeter is the distance around, or 4x.

4x = 48
x = 12

The length of each side is 12m.

The area of the square is the side length times the side length, or x*x. 

12*12 = 144

The area of the square is 144 square meters.

We are told that the difference between the area of the square and the rectangle is 4 square 
meters. Then the area of the rectangle must be 4 more or 4 less. 

For now, assume that the rectangle has the lesser area, or 140 square meters.

The formula for the area of a rectangle is length times width, or L*W. We have
L*W = 140.

We also know that the perimeter of the rectangle is 48. We have
2L + 2W = 48

We have two equations with two variables. Solve for L and W.
Solve the first equation for L.
L*W = 140
L = 140/W

Substitute 140/W for L in the second equation.
2L + 2W = 48
2(140/W) + 2W = 48

Solve for W.
280/W + 2W = 48

Subtract 2W from both sides.
280/W = -2W + 48

Multiply both sides by W.
280 = -2W^2 + 48W

This is a quadratic equation. Solve by factoring.
2W^2 - 48W + 280 = 0

Divide each term by 2.
W^2 - 24W + 140 = 0

(W - 14)(W - 10) = 0

W = 14 or W = 10

The width of the rectangle is 10m or 14m.

If the width of the rectangle is 10m, then its length is 14m because 
2L + 2W  = 2(10) + 2(14) = 48

The dimensions of the rectangle are 10m by 14m.

You can work out the dimensions of the rectangle if it has the larger area. Then your
equations will be

L*W = 148
2L + 2W = 48

I'll leave that to you.

Hope this helps! Feel free to email if you have any questions about the solution.

Good luck with your math,

Mrs. F
math.in.the.vortex@gmail.com